A concrete fence 7 ft. high and 210ft. long is to be painted on both sides.
a. what is the total area of the fence to be painted?
b. A galloon of a certain type of a paint will cover only 200 sq. ft. for the first coat,but on the second coat a galloon of the same paint will cover 300 sq. ft. How many galloons of paint should be bought if the fence is to be given two coats of paint?
That's my question. I really don't know what formula will I use.
kindly explain it for me. :)
thank you, godbless :))

Answer by Cromlix(4381)

Total Area = 2 * height * length
Total Area = 2 * 7ft * 210ft = 2940ft^2
First coat of paint:
2940/200 = 14.7 gallons
Second coat of paint:
2940/300 = 9.8 gallons
Total quantity = 14.7 + 9.8
= 24.5 gallons of paint.
Hope this helps.
:-)